OpenStudy (anonymous):

check my answer to -6/7 * 5 * (-14/27)? -6/7 * 5 * (-14/27) -6/7 *5/1*(-14/27) -30/7 * -70/27= 210/189

OpenStudy (anonymous):

-30/7 *(-14/27) continue from here...

OpenStudy (anonymous):

no that's not correct .... take a calculator and you'll see !

OpenStudy (anonymous):

This is my husband's account, I should probably make one of my own, but I am HORRIBLE at math and am just making sure that I am doing everything right because I can't fail this math class a second time. Not to mention how expensive these classes cost each time!

OpenStudy (anonymous):

My mistake, it's 20/9

OpenStudy (anonymous):

-6/7 *5/1*(-14/27) here you made the mistake: you multiplied like: \[\LARGE \frac{-6\cdot 5}{7} \cdot \frac{-14\cdot 5}{27}\] but in reality it should be like: \[\LARGE \frac{-6\cdot 5}{7} \cdot \frac{-14}{27}\] or.. \[\LARGE \frac{-6}{7} \cdot \frac{-14\cdot 5}{27}\]

OpenStudy (alexwee123):